Need assistance with parts of section C, please

Analyzing Financial Statement Effects of Bond Redemption

Holthausen Corporation issued $100,000 of 11%, 20-year bonds at 108 on January 1, 2010. Interest is payable semiannually on June 30 and December 31. Through January 1, 2016, Holthausen amortized $1,048 of the bond premium. On January 1, 2016, Holthausen retires the bonds at 103.

Required
a. Prepare journal entries to record the issue and retirement of these bonds.

General Journal
Date Description Debit Credit
01/01/10 CASH 108000

0

Bonds payable 0 100000
Bond premium 0

8000

01/01/16 Bonds payable 100000

0

Bond premium 6952

0

Cash 0

103000

Gain on retirement 0

3952

b. Post the journal entries from part a to their respective T-accounts.

Cash (A)
01/01/10 108000

0

01/01/16 0

103000

Bonds Payable (L)
01/01/10 0

100000

01/01/16 100000 0
Gain on Retirement of Bonds (R)
01/01/10 0

0

01/01/16 0 3952
Bond Premium (L)
01/01/10 0

8000

01/01/16 6952 0

c. Record each of the transactions from part a in the financial statement effects template. Remember to use negative signs, when appropriate.

Balance Sheet Income Statement
Transaction Cash Asset + Noncash
Asset
= Liabilities + Contrib.
Capital
+ Earned
Capital
Revenues - Expenses = Net Income
1/1/10 Issue bonds at a premium 108000 + 0 = Answer + 0 + Answer 0 - 0 = 0
1/1/16 Retire bonds issued on 1/1/10 (103000) + 0 = Answer + 0 + Answer Answer - 0 = Answer
Answer #1

c.

Balance Sheet Income Statement
Transaction Cash Asset + Noncash Asset = Liabilities + Contrib. Capital + Earned Capital Revenues - Expenses = Net Income
1/1/10 Issue bonds at a premium 108000 + 0 = 108000 + 0 + 0 0 - 0 = 0
1/1/16 Retire bonds issued on 1/1/10 -103000 + 0 = -106952 + 0 + 3952 3952 - 0 = 3952